### v1

# Can LLMs Generate Diverse Molecules? Towards Alignment with Structural Diversity

Abstract
Recent advancements in large language models (LLMs) have demonstrated impressive performance in generating molecular structures as drug candidates, which offers significant potential to accelerate drug discovery. However, the current LLMs overlook a critical requirement for drug discovery: proposing a diverse set of molecules. This diversity is essential for improving the chances of finding a viable drug, as it provides alternative molecules that may succeed where others fail in wet-lab or clinical validations. Despite such a need for diversity, the LLMs often output structurally similar molecules from a given prompt. While decoding schemes like beam search may enhance textual diversity, this often does not align with molecular structural diversity. In response, we propose a new method for fine-tuning molecular generative LLMs to autoregressively generate a set of structurally diverse molecules, where each molecule is generated by conditioning on the previously generated molecules. Our approach consists of two stages: (1) supervised fine-tuning to adapt LLMs to autoregressively generate molecules in a sequence and (2) reinforcement learning to maximize structural diversity within the generated molecules. Our experiments show that (1) our fine-tuning approach enables the LLMs to better discover diverse molecules compared to existing decoding schemes and (2) our fine-tuned model outperforms other representative LLMs in generating diverse molecules, including the ones fine-tuned on chemical domains.
1 Introduction
Figure 1: LLMs generate a molecule with SMILES representation.
Recent advances in large language models (LLMs) have demonstrated remarkable potential to accelerate scientific discovery by leveraging their language processing capabilities. This progress has been particularly impactful for candidate design problems such as drug discovery (Pei et al., 2024), protein sequence design (Zhuo et al., 2024), and material design (Gruver et al., 2024). In particular, with expansive biomolecular datasets and molecular string representations, e.g., SMILES (Weininger, 1988) or SELFIES (Krenn et al., 2020), LLMs have demonstrated impressive abilities to generate molecules from textual descriptions, e.g., molecular properties (Edwards et al., 2022; Ye et al., 2023; Pei et al., 2024), as illustrated in Figure 1.
However, current LLM-based molecular generation approaches (Edwards et al., 2022; Ye et al., 2023; Pei et al., 2024) often overlook a critical requirement for drug discovery: proposing a diverse set of molecules. In computer-aided drug discovery pipelines, identifying a single molecule with a desired property does not guarantee success in real-world pipelines that require additional cell-based studies and clinical trials (Vamathevan et al., 2019). Therefore, drug discovery requires a collection of structurally diverse molecules, as illustrated in Figure 2. The generation of structurally diverse molecules increases the chances of finding a viable drug candidate (Xie et al., 2023), as different molecules may succeed where others fail. This diversity is essential to enhance the robustness and success of the drug discovery process (Krantz, 1998; Hong et al., 2020; Sadybekov & Katritch, 2023).
Figure 2: The compute-aided drug design. A collection of structurally diverse molecules is required to increase the chance of identifying viable drug candidates in the real world.
In response, we explore the use of LLMs for diverse molecular generation. We begin by identifying the limitations of recent LLMs (Ye et al., 2023; OpenAI, 2023) and decoding schemes (Vijayakumar et al., 2018; Su et al., 2022) in generating diverse molecules. Then, we present a new method for fine-tuning LLMs to generate diverse molecules. Our approach can be broadly applied to other LLM-based candidate design problems, e.g., computer-aided design (Wu et al., 2023a).
Existing LLMs have limitations in generating diverse molecules. To obtain diverse molecules, one may consider querying the recently developed generalist LLMs, e.g., Llama (Touvron et al., 2023) or ChatGPT (OpenAI, 2023). However, our empirical observation in Figure 3(a) reveals that even the most recent models produce structurally identical or highly similar molecules from the given prompt. [^1] ChatGPT-4o (OpenAI, 2023) generates different SMILES strings that map to an identical molecule. This observation aligns with previous observations that have shown LLMs may fail to generate diverse outputs (Kirk et al., 2024) for general text-based domains.
Decoding schemes for diversified generation do not align with molecular diversity. We also acknowledge the existence of decoding schemes, e.g., diverse beam search (Vijayakumar et al., 2018) or contrastive beam search (Su et al., 2022), which have been proposed to improve the diversity of output sequences generated by LLMs. However, these decoding schemes are limited to improving the textual diversity which often does not correspond to molecular structural diversity, e.g., there exist many SMILES or SELFIES strings that correspond to the same molecule, as illustrated in Figure 3(b).
(a) Existing LLMs (Ye et al., 2023; OpenAI, 2023) lack the ability to generate a diverse set of molecules.
(b) (Left) Diverse output sequences (SMILES) induce the same molecular structures. (Right) Improved textual diversity via diverse beam search does not enhance molecular diversity in the experiments (Section 4.1).
Figure 3: Existing works on LLMs fail to generate diverse molecules. The existing decoding schemes (Vijayakumar et al., 2018) for diverse sequence generation and LLMs for chemical tasks fail to capture the molecular diversity, and may induce structurally identical molecules.
Our approach. We repurpose existing molecular generative LLMs to autoregressively generate a diverse set of molecules from a single prompt. By enabling the LLMs to generate a new molecule conditioned on previously generated molecules, we expect the LLMs to learn to enhance the structural diversity between the generated molecules. To this end, we propose a two-stage approach to fine-tune LLMs: (a) a supervised fine-tuning stage to repurpose LLMs to autoregressively generate a sequence of multiple molecules and (b) a reinforcement learning stage to maximize the molecular structural diversity. Note that both stages do not require external diverse molecular datasets and are purely based on self-improvement procedures, where the LLMs train on the samples generated by themselves.
In the supervised training stage, we train the LLMs to autoregressively generate a set of molecules in a single sequence. The training dataset, i.e., a set of molecules, is collected from the LLMs themselves through iterative sampling, and then filtered to enhance the quality, e.g., removing invalid molecules. However, this stage does not necessarily incorporate molecular diversity, as the training may not involve sufficiently distinct molecules (e.g., limitations in Figure 3(b)). To tackle this, we subsequently apply reinforcement learning with exploration towards discovering diverse molecules.
Next, in the reinforcement learning stage, we train LLMs to maximize the diversity of molecules within a generated sequence. However, for our task, conventional sequence-wise reinforcement learning (Ouyang et al., 2022) suffers from the credit assignment problem (Zhou et al., 2024): the challenges in identifying and promoting the generation of molecules responsible for increasing diversity, among a larger set of molecules in the sequence. To resolve this issue, we solve multi-stage molecule generation problems for a sequence of molecules, where the generation of each molecule aims to maximize the diversity with respect to the previously generated molecules. We train LLMs to maximize the associated rewards using proximal policy optimization (Schulman et al., 2017).
We compare our method with the decoding schemes for diversified generation (Vijayakumar et al., 2016; Su et al., 2022) and other representative LLMs, including chemical-task specialists (Edwards et al., 2022; Christofidellis et al., 2023; Pei et al., 2023; 2024), fine-tuned generalists on chemical domains (Fang et al., 2024; Yu et al., 2024), and the ChatGPT series (OpenAI, 2023; 2024). We observe that (1) our fine-tuning approach enables LLMs to better discover diverse molecules compared to existing decoding schemes and (2) our fine-tuned LLM outperforms other recent LLMs.
To conclude, our contributions can be summarized as follows:
  • We are the first to explore the use of LLMs for generating diverse molecules.
  • We first propose a fine-tuning approach for LLMs to generate diverse solutions, which presents a new direction distinct from existing approaches focused on the decoding scheme.
  • Experimentally, our method outperforms the baselines in generating diverse molecular structures.
2 Related work
Large language models (LLMs) for molecular generation. Recent advancements in LLMs have shown increasing promise in scientific applications, especially for molecular generation (Edwards et al., 2022; Pei et al., 2023; Fang et al., 2024; Pei et al., 2024). First, Edwards et al. (2022) proposed MolT5 which translates between SMILES (Weininger, 1988) and molecular text descriptions. Text+Chem T5 (Christofidellis et al., 2023) is a model pre-trained on both the chemical and natural language domains. Next, BioT5 (Pei et al., 2023) considers T5 models pre-trained on datasets including bio-text, protein sequences, and molecules. Additionally, Ye et al. (2023), Fang et al. (2024), and Yu et al. (2024) fine-tuned generalist LLMs, e.g., Llama (Touvron et al., 2023), through biological instructions, molecular modifications, and large-scale molecular datasets, respectively.
Decoding schemes for generating diverse output sequences. To generate diverse and high-quality solution candidates from LLMs, existing literature on LLMs has studied improving decoding schemes. To acquire multiple distinct sequences with high likelihoods, one can consider employing beam search, which jointly decodes multiple distinct outputs (Och, 2003). To enhance diversity between sequences, Vijayakumar et al. (2016; 2018) incorporated token-wise differences between generated sequences in the beam search. Furthermore, Su et al. (2022) considered the contrast between the candidate sequences. In addition, Holtzman et al. (2020) proposed nucleus sampling, which enhances random sampling by balancing the quality and the diversity.
Reinforcement learning (RL) for fine-tuning LLMs. RL has been effectively applied to fine-tune LLMs, aligning them with desired behaviors expressed through reward signals. One notable example is RL from human feedback to align LLMs with human preference (Ouyang et al., 2022). In addition, there has been a surge in research on devising RL for LLMs as well, such as addressing multi-turn settings (Shani et al., 2024) and incorporating multiple fine-grained reward signals (Wu et al., 2023b). For molecular generation, Ghugare et al. (2024) proposed RL-based fine-tuning to generate a molecule satisfying target properties. However, to the best of our knowledge, there exist no prior RL-based approaches that aim to increase the diversity of LLM-generated outputs.
(a) Supervised fine-tuning: enables generating a sequence of multiple molecules.
(b) Reinforcement learning: maximizes the reward of generated molecules.
(c) Reward computation: captures the structural diversity.
Figure 4: Illustration of proposed fine-tuning approaches. We consider two stages for fine-tuning LLMs: a supervised fine-tuning Figure 4(a) and a reinforcement learning Figure 4(b). The prompts are simplified for explanatory purposes, and the actual prompts are provided in Appendix B.
3 Method
In this section, we present our method for fine-tuning LLMs to generate diverse molecules. Specifically, we consider applying fine-tuning to existing molecular generative LLMs that produce molecular representations such as SMILES or SELFIES. Importantly, our approach is versatile and can be broadly applied to other domains, e.g., protein sequence (Zhuo et al., 2024) or computer-aided design (Wu et al., 2023a). Furthermore, it leverages self-improvement techniques and does not require additional datasets containing diverse molecules.
Overview. Our goal is to generate a sequence of structurally diverse molecules from a given prompt by producing them in a single concatenated output. To achieve this, we fine-tune the LLMs in two stages: (a) a supervised fine-tuning phase that repurposes the LLMs to generate a sequence of molecules rather than a single one, and (b) a reinforcement learning phase aimed at further enhancing the structural diversity among the generated molecules.
Task details. In detail, we consider generating molecules from a prompt $p_{\text{desc}}$, where the prompt describes a molecular property that the generated molecules should possess. Here, we aim to generate diverse molecules that satisfy the given description $p_{\text{desc}}$. We let $\mathcal{P}$ denote the prompts used for training, which does not intersect with the ones used for evaluation.
Algorithm 1 Supervised fine-tuning
1:  Initialize $\pi_{\text{SFT}}$ with $\pi_{\text{pre}}$
2:  repeat
3:     Sample prompt $p_{\text{desc}}\sim\mathcal{P}$
4:     Sample $\{m_{i}\}_{i=1}^{T}$ from $\pi_{\text{pre}}(m\;|\;p_{\text{desc}})$
5:     Update $\{m_{i}\}^{K}_{i=1}\leftarrow\text{Filter}(\{m_{i}\}^{T}_{i=1})$
6:     Maximize Equation 1 with $\{m_{i}\}^{K}_{i=1}$
7:  until Converged
8:  Output: fine-tuned $\pi_{\text{SFT}}$
3.1 Supervised fine-tuning
We first describe our supervised fine-tuning process for repurposing the pre-trained LLMs to autoregressively generate multiple molecules in a sequence. This involves collecting a dataset of molecules from a pre-trained LLM $\pi_{\text{pre}}$, and then fine-tuning the LLM $\pi_{\text{SFT}}$ on the collected dataset. We describe the process in Figure 4(a) and Algorithm 1.
Dataset collection. The dataset collection and supervised training processes are conducted with a set of training prompts $\mathcal{P}$. Initially, the pre-trained LLM $\pi_{\text{pre}}$ produces a set of molecules by iterative sampling molecules for a given prompt $p_{\text{desc}}\in\mathcal{P}$ as follows:
$m_{i}\sim\pi_{\text{pre}}\left(m_{i}\middle|p_{\text{desc}}\right)\quad\text{for}\;i=1,\ldots,T,$
where $m_{i}$ denotes the string representation of the molecule. In practice, we employ beam search to collect the set of molecules $\{m_{i}\}^{T}_{i=1}$. Then, we filter out the invalid string representations, duplicate molecules, and molecules that do not satisfy the given prompt $p_{\text{desc}}$. This results in reducing the set of molecules from $\{m_{i}\}^{T}_{i=1}$ to $\{m_{i}\}^{K}_{i=1}$. The details are described in Appendix A.
Supervised training. Given the filtered set of molecules $\{m_{i}\}^{K}_{i=1}$, we train the LLM $\pi_{\text{SFT}}$, which is initialized with $\pi_{\text{pre}}$, to generate them as a single concatenated sequence. We denote this sequence by $\mathcal{M}_{1:K}=m_{1}||\cdots||m_{K}$, where ||||| | denotes the concatenation of the molecular string representations. Specifically, given a modified prompt $p_{\text{desc+div}}$ describing the target property with a request for generating diverse molecules, we train the LLM to maximize the log-likelihood:
$\log\pi_{\text{SFT}}\left(\mathcal{M}_{1:K}\;\middle|\;p_{\text{desc+div}}\right).$ (1)
However, the policy $\pi_{\text{SFT}}$ does not necessarily incorporate a molecular structural diversity, as the set of molecules $\{m_{i}\}^{K}_{i=1}$ collected from $\pi_{\text{pre}}$ may insufficiently involve diverse molecular structures (e.g., due to limitations in Figure 3(b)). To tackle this, we next introduce an online reinforcement learning stage with exploration towards discovering diverse molecules.
3.2 Reinforcement learning
Algorithm 2 Multi-stage RL fine-tuning
1:  Sample $p_{\text{desc}}\sim\mathcal{P}$
2:  Sample $m_{1}\sim\pi_{\text{RL}}(m_{1}\mid p_{\text{desc+div}})$
3:  Update $\pi_{\text{RL}}$ with PPO to maximize $r(m_{1})$
4:  for $k=2,\ldots,K$ do
5:     Sample $m_{k}\sim\pi_{\text{RL}}(m_{k}\mid\mathcal{M}_{1:k-1},p_{\text{desc+div}})$
6:     Update $\pi_{\text{RL}}$ with PPO to maximize $r(m_{k})$
7:  end for
We apply reinforcement learning to maximize the diversity of the generated molecules within a sequence. However, when applied to a sequence of molecules $\mathcal{M}_{1:K}$, conventional sequence-wise reinforcement learning (Ouyang et al., 2022) suffers from the credit assignment problem (Zhou et al., 2024): the challenge in identifying and promoting the generation of molecules responsible for increasing diversity, among a set of molecules $\{m_{i}\}^{K}_{i=1}$. To circumvent this, we introduce a molecule-wise reinforcement learning. [^2] We compare both approaches in Table 4 of Section 4.4.
Specifically, we consider reinforcement learning on a sequence of molecules $\mathcal{M}_{1:K}$ as learning in $K$ individual stages. Each stage corresponds to generating a molecule $m_{k}$ conditioned on a sequence of previously generated molecules $\mathcal{M}_{1:k-1}$. Then, the LLM $\pi_{\text{RL}}$ is trained to maximize the return of each stage, which is defined by the reward of the generated molecule $m_{k}$. The reward is defined as the diversity between the previously generated molecules $\{m_{i}\}_{i=1}^{k-1}$ and the new molecule $m_{k}$. We also incorporate an auxiliary reward to ensure that the molecule $m_{k}$ satisfies the given description $p_{\text{desc}}$. We illustrate our approach in Figure 4(b) and provide the detailed algorithm in Algorithm 2.
Reward. The reward evaluates the generated molecule $m_{k}$ with a diversity reward $r_{\text{div}}(m_{k},\{m_{i}\}^{k-1}_{i=1})$ and a description-matching reward $r_{\text{match}}(m_{k},p_{\text{desc}})$, as follows:
$r(m_{k})=r_{\text{div}}(m_{k},\{m_{i}\}^{k-1}_{i=1})+r_{\text{match}}(m_{k},p_{\text{desc}}),$
where the diversity reward $r_{\text{div}}$ evaluates structural differences between the molecule $m_{k}$ and the previously generated molecules $\{m_{i}\}^{k-1}_{i=1}$ by assessing their true molecular structures. Note that $r_{\text{div}}(m_{1})$ is zero. The description-matching reward $r_{\text{match}}$ evaluates whether the molecule $m_{k}$ satisfies the given description $p_{\text{desc}}$. In a practical implementation, the final action that completes the molecule $m_{k}$ yields the reward $r(m_{k})$. The detailed reward implementation is described in Appendix A.
Policy optimization. We optimize the LLM $\pi_{\text{RL}}$ to maximize the reward using proximal policy optimization (PPO; Schulman et al., 2017; Ouyang et al., 2022). Note that $\pi_{\text{RL}}$ is initialized with $\pi_{\text{SFT}}$. In addition, we combine per-token KL penalty from the supervised fine-tuned model at each token following prior studies (Ouyang et al., 2022).
4 Experiment
In this section, we validate our supervised fine-tuning and reinforcement learning methods for generating diverse molecules, coined Div-SFT and Div-SFT+RL, respectively. In these experiments, one can observe that (1) our fine-tuning approach enables LLMs to better discover diverse molecules compared to decoding schemes for diversified generation and (2) our fine-tuned LLM outperforms other representative LLMs, including generalist and specialist models for chemical tasks.
4.1 Comparison with decoding schemes
In this experiment, we show that our method enables LLMs to better generate diverse molecules compared to the existing decoding schemes for diverse sequence generation. To validate the consistent improvement, we implement our fine-tuning method and decoding schemes on two models specialized in the chemical domain: BioT5+ (Pei et al., 2024) and MolT5 (Edwards et al., 2022).
(a) Comparison with decoding schemes on BioT5+
(b) Comparison with decoding schemes on MolT5
Figure 5: Comparison with decoding schemes. NCircles represents both quality and diversity-related metric. Top 10 and Accepted & unique represent quality-related metrics. IntDiv. represents a diversity-related metric. Our method generates more diverse and high-quality molecules compared to the baselines. Notably, our method makes a larger gap over the baselines on metrics related to capturing both quality and diversity, i.e., NCircles.
Tasks. We consider description-guided molecule generation using the ChEBI-20 dataset, which includes training and test sets (Edwards et al., 2021). Note that the training dataset has also been used to pre-train the base LLMs, BioT5+ and MolT5, i.e., not an external dataset. The test dataset is unobserved during both the pre-training and fine-tuning phases. Each data point consists of a pair of a description and a ground-truth example molecule that satisfies the description.
In our experiments, we consider generating $50$ molecules for each description in the ChEBI-20 test dataset involving $3300$ molecular descriptions. We then evaluate both the structural diversity of the generated molecules and how well they satisfy the given descriptions.
Metrics. To measure the similarity between two molecules, we use Tanimoto similarity (Bajusz et al., 2015), which based on distance of molecular structures from fingerprints. To be specific, we consider the following metrics for generated molecules:
  • The number of accepted and unique molecules (Accepted & unique): This metric measures the number of valid and unique molecules satisfying the given description. [^3] Following the evaluation metric of prior studies (Edwards et al., 2022; Pei et al., 2024), we evaluate the molecule as satisfying the given description when its BLEU score between a ground-truth sample is high ($>0.7$).
  • The number of circles (NCircles; Xie et al., 2023): This metric considers both the quality and diversity of generated molecules. For the accepted molecules satisfying the given description, this metric counts the number of maximum distinct clusters with Tanimoto similarity (Bajusz et al., 2015) below a threshold $h$, i.e., indicating chemical space coverage (illustrated in Figure 6).
  • Internal diversity (IntDiv.; Polykovskiy et al., 2020): This metric is the complement of the average of pair-wise Tanimoto similarities between the accepted molecules satisfying the given description.
  • Average of top 10 scores (Top 10): This metric measures the quality of the generated molecules by averaging the BLEU scores (Papineni et al., 2002) of the top unique $10$ molecules yielding high BLEU scores between the ground-truth example.
(a) BS
(22,11)
(b) Diverse BS
(7,4)
(c) Contrastive BS
(26, 11)
(d) Div-SFT+RL
(29,28)
Figure 6: The distribution of (Accepted & unique, NCirclesh=0.65) obtained by each method. The molecules are generated from the description in Table 13. The Fruchterman-Reingold force-directed algorithm (Fruchterman & Reingold, 1991) embeds each molecule in a 2-dimensional space based on Tanimoto similarity. Here, a low distance yields a high Tanimoto similarity. Each dot, shaded circle, and edge represent accepted molecules, distinct clusters (used to measure NCircles), and a pair with similarity above $0.65$, respectively. Our method better captures the molecular diversity.
Table 1: Visualization of the generated molecules with their diversity. The eight molecules are generated with the description: “The molecule is a primary aliphatic ammonium ion which is obtained from streptothricin F by protonation of the guanidino and amino groups. It has a role as an antimicrobial agent. It is a guanidinium ion and a primary aliphatic ammonium ion. It is a conjugate acid of a streptothricin F’.” The generated molecule with blue line indicates the accepted molecule.
Method (IntDiv.) Example of outputs
BS (0.51) [Uncaptioned image]
Diverse BS (0.28) [Uncaptioned image]
Contrastive BS (0.53) [Uncaptioned image]
Div-SFT+RL (0.69) [Uncaptioned image]
Implementations. For supervised fine-tuning, we collect a hundred molecules for each molecular description $p_{\text{desc}}$ in ChEBI-20 training sets. Then, they are filtered to remove invalid molecules, duplicated molecules, and unaccepted molecules. The description-matching reward $r_{\text{match}}(m_{k})$ is defined as a BLEU score between a ground-truth example. The diversity reward $r_{\text{div}}(m_{k},\{m_{i}\}^{k-1}_{i=1})$ is defined as the complement of the maximum Tanimoto similarity between generated molecules $\{m_{i}\}^{k-1}_{i=1}$. The detailed implementations and hyper-parameters follow Appendix A.
Baselines. We compare our fine-tuning approach with various decoding schemes. We consider the random sampling with different temperatures $\{0.7,1.0,1.5\}$ and nucleus sampling (Holtzman et al., 2020). We also consider beam search (BS) and the variants of BS to promote sequence-level diversity between the generated samples: diverse BS (Vijayakumar et al., 2018) and contrastive BS (Su et al., 2022). The detailed settings are described in Appendix B.
Results. We present the main results on BioT5+ and MolT5 in Figure 5(a) and Figure 5(b), respectively. One can see that our approach, i.e., Div-SFT+RL, shows superior performance compared to the considered baselines. Especially, it is worth noting that our approach makes significant improvements in NCircles metrics that require generating diverse molecules satisfying the given description.
We also provide the qualitative results in Figure 6 and Table 13 with sampled molecules. One can see that variants of beam search for diverse sentence generation, i.e., diverse BS and contrastive BS, do not enhance molecular structural diversity. In contrast, our approach demonstrates significantly higher molecular structural diversity. In Appendix C, we provide additional examples of molecules generated by the considered baselines and our method.
Table 2: Comparison with existing LLMs on description-based molecule generation. Our fine-tuned model shows superior performance compared to the considered baselines.
Method Accepted & Unique NCirclesh=0.75 NCirclesh=0.65 IntDiv. Top 10
Chemical-task Specialist LLMs
MolT5 3.12 1.85 1.59 0.29 0.41
Text+Chem T5 13.78 4.82 3.29 0.23 0.63
BioT5 15.58 8.45 6.07 0.33 0.69
BioT5+ 16.03 7.93 6.16 0.31 0.63
Generalist LLMs
Mol-instructions 0.06 0.03 0.03 0.02 0.01
LlaSMol 17.68 6.73 4.94 0.26 0.68
GPT-3.5-turbo 3.87 2.83 2.49 0.20 0.41
GPT-4o 5.74 4.09 3.53 0.22 0.48
o1-preview 6.92 3.46 2.36 0.13 0.31
BioT5++ours 21.98 16.98 14.35 0.45 0.74
4.2 Comparison with existing LLMs
Here, we compare our fine-tuned BioT5+, presented in Figure 5, with various recent LLMs, including chemical-task specialists, generalists, and fine-tuned generalist LLMs on chemical domains. The purpose of this experiment is to highlight the limitations of existing LLMs in capturing molecular diversity, whereas our fine-tuned model successfully captures molecular diversity during generation. The tasks and metrics are the same as settings in Section 4.1. We use the first $500$ molecular descriptions in the ChEBI-20 test dataset for evaluation.
Baselines. We compare our fine-tuned model with various existing LLMs. We consider four LLMs specialized for chemical tasks: Text+Chem T5 (Christofidellis et al., 2023), MolT5, BioT5 (Pei et al., 2023), and BioT5+. Next, we consider two generalist LLMs: Mol-instructions (Fang et al., 2024) fine-tuned from Llama-7B (Touvron et al., 2023), and LlaSMol (Yu et al., 2024) fine-tuned from Mistral-7B (AI, 2023), based on description-based molecular generation tasks. For each baseline, we report the best results (highest NCirclesh=0.65) obtained using either random sampling, beam search, diverse beam search, or contrastive beam search. We also consider three strong API-based generalist LLMs: GPT-3.5, GPT-4o, and o1-preview (OpenAI, 2023; 2024). [^4] We use gpt-3.5-turbo-0125 and gpt-4o-2024-08-06. The detailed experimental settings and prompts are described in Appendix B.
Results. We present the results in Table 2. One can see that our fine-tuned model shows superior performance compared to the considered baselines in discovering diverse and high-quality molecules. Furthermore, it is noteworthy that most existing LLMs yield low NCircles with respect to the number of generated accepted and unique molecules, while our method yields relatively high NCircles.
4.3 Fine-tuning generalist LLMs
We further validate whether our fine-tuning method improves the generalist LLMs as well in terms of the diversity of generated molecules. Here, as a base LLM for implementing our method, we consider DrugAssist (Ye et al., 2023) which is fine-tuned from the Llama-7B. As baselines, we apply random sampling and contrastive beam search to DrugAssist.
Tasks. In this experiment, we consider prompts based on the four quantitative molecule properties: hydrogen bond (HB) donors, HB acceptors, Bertz complexity (Bertz, 1981), and quantitative estimate of drug-likeness (QED) (Bickerton et al., 2012). The goal of this task is to generate diverse molecules that satisfy the properties described in the given prompt.
Implementations. We apply our supervised fine-tuning and reinforcement learning to enhance DrugAssist. For supervised fine-tuning, we collect multiple molecules for prompts about three properties: HB donors, HB acceptors, and Bertz complexity. The prompts about QED are excluded from training but included in the evaluation to assess generalization to unseen properties. The sampled molecules are filtered to remove invalid molecules, duplicated molecules, and molecules that do not satisfy the given properties. The property-match reward $r_{\text{match}}(m_{k})$ yields a non-zero value when satisfying the given property, and the diversity reward $r_{\text{div}}(m_{k},\{m_{i}\}^{k-1}_{i=1})$ is defined as same as Section 4.1. The detailed implementations and hyper-parameters follow Appendix A.
(a) HB donors
(b) HB acceptors
(c) Bertz complexity
(d) QED
Figure 7: Experiments on DrugAssist. Our method consistently improves the performance for generating diverse and high-quality molecules when implemented on generalist LLMs.
Results. We present the results in Figure 7. One can see that our approach consistently improves the performance in generating diverse and high-quality molecules when implemented on generalist LLMs. Note that our approach consistently demonstrates superior performance for the unseen prompt, i.e., the prompt about QED in Figure 7(d).
Table 3: Experiments with the large number of samples. Our method discovers more diverse molecules with respect to the (1) the number of generations and (2) time costs.
Methodnum. of ${}_{\text{num. of generations}}$ ${\text{BS}_{300}}$ ${\text{BS}_{400}}$ ${\text{BS}_{500}}$ ${\text{Ours}_{85}}$ ${\text{Ours}_{120}}$ ${\text{Ours}_{155}}$
NCirclesh=0.65 18.6 20.4 21.3 20.4 23.6 25.6
Time (sec.) 323 452 585 65 86 107
4.4 Ablation studies
Table 4: Comparison with variants of implementations. Applying multi-stage reinforcement learning shows superior performance compared to (1) supervised fine-tuning with hard filtering and (2) single-stage reinforcement learning.
Method Accepted & Unique NCirclesh=0.75 NCirclesh=0.65 IntDiv. Top 10
[…]
Div-SFT+RL 21.98 16.98 14.35 0.45 0.74
Large number of samples vs. performance. We also analyze how well our method discovers diverse molecules with respect to the number of generations. In this experiment, we extend beyond the settings in Section 4.1. We consider our method to generate $85$, $120$, and $155$ molecules with a single NVIDIA A100 SXM4 40GB GPU. Next, we consider beam search with beam sizes of $300$, $400$, and $500$, respectively. The beam search is implemented with four NVIDIA A100 SXM4 40GB GPUs due to the memory limitation of a single GPU. Note that this experiment uses $250$ molecular descriptions in the ChEBI-20 test set. We present the results in Table 3. One can see that our method (1) discovers more diverse molecules with respect to the total number of generations and (2) exhibits further performance improvements as the number of generations increases.
Time costs vs. performance. In addition, we also analyze how well our method discovers diverse molecules with respect to the time costs. In Table 3, we present the time costs for each method. One can see that our method discovers more diverse molecules in a practical time compared to the decoding schemes for diverse sequence generation.
SFT with hard filtering vs. RL. To train LLMs to generate diverse molecules, one may also consider supervised fine-tuning on hard-filtered datasets, e.g., using a set of molecules filtered by similarity, as an alternative to reinforcement learning. In this experiment, we additionally perform supervised fine-tuning with distinct molecules, where each pair of molecules yields a Tanimoto similarity below $0.65$. The results are presented in Table 4. One can see that while it yields relatively high NCircles with respect to the number of accepted unique molecules, the overall performance is worse compared to the performance of reinforcement learning.
Single-stage vs. multi-stage RL. As mentioned in Section 3.2, we consider a multi-stage setting for generating multiple molecules. However, one may also consider a single-stage setting, where the return of a generated sequence is defined as the sum of the rewards from multiple generated molecules. In Table 4, we compare both approaches. One can see that the multi-stage setting significantly outperforms the single-stage setting. We hypothesize that this result stems from credit assignment issues in the single-stage setting. Namely, the single-stage setting lacks signals to capture molecule-wise impacts on diversity among a large set of molecules and fails to promote the generation of molecules responsible for increasing diversity.
5 Conclusion
In this paper, we identify the limitations of large language models (LLMs) for generating diverse molecules. Then, we present new supervised fine-tuning and reinforcement learning methods to adapt existing molecular generative LLMs to generate a diverse set of molecules. Experiments show that our approach enables LLMs to better discover diverse molecules compared to the existing approaches. An interesting avenue for future work is to reduce the space and time complexity in generating the sequence of molecules, for example, based on set encoding techniques (Zaheer et al., 2017).
Appendix A Detailed implementations and training
A.1 Supervised fine-tuning
Dataset Collection. To fine-tune BioT5+ and MolT5 (Sections 4.1 and 4.2), we collect $T=100$ molecules using beam search for each training molecular description in the ChEBI-20 training dataset. Note that this dataset has been considered in the original BioT5+ and MolT5.
[…]
Supervised learning. We consider four NVIDIA A100 GPUs for supervised fine-tuning.
  • For the supervised fine-tuning of BioT5+ and MolT5 (Sections 4.1 and 4.2), we consider $80$ epochs, $8$-batch size, $5\textrm{e}-4$ learning rate, $0.05$ warm-up ratio, and apply a cosine learning scheduler. The maximum sequence length in supervised training is limited to $2560$ due to memory limitations.
  • For the supervised fine-tuning of DrugAssist (Section 4.3), we consider $80$ epochs, $4$-batch size, $3\textrm{e}-5$ learning rate, $0.05$ warm-up ratio, and apply a cosine learning scheduler. The maximum sequence length in supervised training is limited to $1024$ due to memory limitations. We also apply LoRA (Hu et al., 2022), where the rank and alpha are $64$ and $128$, respectively.
A.2 Reinforcement learning
Reward Design. In experiments with BioT5+ and MolT5 on the ChEBI-20 dataset (Sections 4.1 and 4.2), we define the description-matching reward using the BLEU score as follows:
$r_{\text{match}}(m_{k},p_{\text{desc}})=\text{BLEU}(m_{p_{\text{mol}}},m_{k})^{\alpha},$ (2)
where $m_{p_{\text{mol}}}$ is a ground-truth molecule satisfying the given description. The ChEBI-20 dataset involves a set of pairs $(p_{\text{desc}},m_{p_{\text{mol}}})$. Note that $\alpha$ is a hyper-parameter.
For experiments with DrugAssist (Section 4.3), the property-matching reward yields $1$ if the molecule satisfies the quantitative properties described in $p_{\text{desc}}$, e.g., HB donors, HB acceptors, and Bertz complexity, and 00 otherwise. The properties are evaluated with the RDKit package (Landrum, 2016).
Next, the diversity reward, $r_{\text{div}}$, is defined to consider molecular structural diversity as follows:
$r_{\text{div}}(m_{k},\{m_{i}\}^{k-1}_{i=1})=1-\text{max}_{m\in\{m_{i}\}^{k-1}_{i=1}}T(m_{k},m)^{\beta},$ (3)
where $T(m_{i},m_{j})$ measures the Tanimoto similarity between $m_{i}$ and $m_{j}$ by assessing their true molecular structures, i.e., fingerprints. This metric yields a value between 00 and $1$ and can be obtained using the RDKit package. Note that $\beta$ is a hyper-parameter.
Policy optimization. We consider four NVIDIA A100 SXM4 40GB for reinforcement learning implemented with proximal policy optimization.
  • For the reinforcement learning of BioT5+ and MolT5 (Sections 4.1 and 4.2), we consider $200$ PPO iterations, $8$ mini-batch size, $128$ batch size, and $5\textrm{e}-5$ learning rate. We also consider $0.01$ KL penalty. Note that $\alpha$ and $\beta$ in Equations 2 and 3 are $0.5$ and $2.0$, respectively. The reward signal is amplified by multiplying by a value of $8.0$. The maximum sequence length in reinforcement learning is limited to $2560$ due to memory limitations. Here, we also apply LoRA (Hu et al., 2022) where the rank and alpha are $16$ and $32$, respectively. We save the model every $25$ PPO iteration and select the model yielding the highest rewards for the training prompts.
  • For the reinforcement learning of DrugAssist (Section 4.3), we consider $200$ PPO iterations, $4$ mini-batch size, $64$ batch size, and $3\textrm{e}-6$ learning rate. We also consider $0.1$ KL penalty. Note that $\beta$ in Equation 3 is $2.0$. The reward signal is amplified by multiplying by a value of $4.0$. The maximum sequence length in reinforcement learning is limited to $1280$. We also apply LoRA (Hu et al., 2022) where the rank and alpha are $64$ and $128$, respectively. We save the model every $25$ PPO iteration and select the model yielding the highest rewards for the training prompts.
Appendix B Detailed experimental settings
Comparison with decoding schemes (Section 4.1). In this experiment, we first consider random sampling with different temperatures $\{0.7,1.0,1.5\}$. For the other decoding schemes, we consider conventional configurations: nucleus sampling with top-p $0.8$, beam search, diverse beam search with a diversity penalty of $0.5$, and contrastive beam search with a penalty alpha of $0.5$. We apply greedy decoding for our approach. The prompts for BioT5+ are described in Table 5. The prompts for MolT5 are considered as a molecular description without any additional comments.
Table 5: Prompts for BioT5+ (Pei et al., 2024).
Prompt Contents
$p_{\text{desc}}$ “Definition: You are given a molecule description in English. Your job is to generate the molecule SELFIES that fits the description. Now complete the following example - Input: <molecular description> Output: ”
$p_{\text{desc+div}}$ (fine-tuning) “Definition: You are given a molecule description in English. Your job is to generate the molecule SELFIES that fits the description. Now provide a set of molecules - Input: <molecular description> Output: ”
Comparison with LLMs (Section 4.2). In this experiment, we consider generalist LLMs. For Mol-instructions, and LlasMol, we apply random sampling, beam search, diverse beam search with a diversity penalty of $0.5$, and contrastive beam search with a penalty alpha of $0.5$. The prompts are described in Table 6. For ChatGPT, we apply either $50$ random sampling with $p_{\text{desc}}$ or greedy decoding with $p_{\text{desc+div}}$, where both prompts are described in Table 7. The prompt design is based on (Li et al., 2024). For o1, we use the prompt $p_{\text{desc+div}}$ described in Table 7. We report the test score using the decoding scheme or the prompt that yields the highest NCirclesh=0.65 for $500$ evaluation molecular description.
Table 6: Prompts for generalist LLMs (Fang et al., 2024; Taylor et al., 2022).
Method $p_{\text{desc}}$
Mol-instructions “Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. ### Instruction: Create a molecule with the structure as the one describe. ### Input: <molecular description> ### Response: ”
LlasMol “Give me a molecule that satisfies the conditions outlined in the description: <molecular description>”
Table 7: Prompts for ChatGPT and o1 (Li et al., 2024).
Input Contents
System $p_{\text{desc}}$ “You are now working as an excellent expert in chemistry and drug discovery. Given the caption of a molecule, your job is to predict the SMILES representation of the molecule. The molecule caption is a sentence that describes the molecule, which mainly describes the molecule’s structures, properties, and production. You can generate the molecule SMILES representation from the caption. Example: Instruction: Given the caption of a molecule, predict the SMILES representation of the molecule. Input: <molecular description>. Your output should be: {"molecule": "<SMILES>"} Your response should only be in the JSON format above; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.”
User $p_{\text{desc}}$ “Input: <molecular description>”
System $p_{\text{desc+div}}$ “You are now working as an excellent expert in chemistry and drug discovery. Given the caption of a molecule, your job is to generate diverse molecules. The molecule caption is a sentence that describes the molecule, which mainly describes the molecule’s structures, properties, and production. You can generate the molecule SMILES representation from the caption. Example: Instruction: Given the caption of a molecule, generate fifty structurally diverse molecules with SMILES representations. Input: <molecular description>. Your output should be: {"molecule1": "<SMILES>"} {"molecule2": "<SMILES>"} {"molecule3": "<SMILES>"} {"molecule50": "<SMILES>"} Your response should only be in the JSON format above; THERE SHOULD BE NO OTHER CONTENT INCLUDED IN YOUR RESPONSE.”
User $p_{\text{desc+div}}$ “Input: <molecular description>”
Fine-tuning generalist LLMs. We synthesize $600$ pairs of training prompts and corresponding sets of molecules. The prompts specify hydrogen bond donors and acceptors ranging from one to four, and a Bertz complexity ranging from 00 to $300$. The sets of molecules are collected by applying beam search on DrugAssist and then perturbed by shuffling their order. We use the prompts described in Table 8. The system prompt follows the default settings from (Ye et al., 2023). For evaluation, we use four prompts specifying three hydrogen bond donors, three hydrogen bond acceptors, a Bertz complexity between $100$ and $200$, and a QED value between $0.4$ and $0.6$. Additionally, as shown in Figure 3, we try to generate diverse molecules by designing prompts without fine-tuning (Table 9). However, these prompts show lower performance compared to applying beam search.
Table 8: Prompts for DrugAssist (Ye et al., 2023).
Prompt Contents
$p_{\text{desc}}$ Hydrogen bond donors and acceptors: “Can you generate a molecule with <value> <property>? Print it in SMILES format.” QED and Bertz complexity: “Can you generate a molecule with <property> below <value1> but at least <value2>? Print it in SMILES format.”
$p_{\text{desc+div}}$ (fine-tuning) Hydrogen bond donors and acceptors: “Can you generate a set of molecules that have <value> <property>? Print each of them in SMILES format.” QED and Bertz complexity: “Can you generate a set of molecules that have <property> below <value1> but at least <value2>? Print each of them in SMILES format.”
Table 9: Prompts for DrugAssist (without fine-tuning, Figure 3).
$p_{\text{desc+div}}$
“Can you generate a set of molecules? Each molecule has <value> <property>. Print each of them in SMILES format.”
“Can you generate a diverse set of molecules? Each molecule has <value> <property>. Print each of them in SMILES format.”
“Can you generate a structurally diverse set of molecules? Each molecule has <value> <property>. Print each of them in SMILES format.”
“Can you generate fifty molecules? Each molecule has <value> <property>. Print each of them in SMILES format.”
“Can you generate fifty diverse molecules? Each molecule has <value> <property>. Print each of them in SMILES format.”
“Can you generate fifty structurally diverse molecules? Each molecule has <value> <property>. Print each of them in SMILES format.”
“Can you generate a set of molecules with <value> <property>? Print each of them in SMILES format.”
“Can you generate a diverse set of molecules with <value> <property>? Print each of them in SMILES format.”
“Can you generate a structurally diverse set of molecules with <value> <property>? Print each of them in SMILES format.”
“Can you generate fifty molecules with <value> <property>? Print each of them in SMILES format.”
“Can you generate fifty diverse molecules with <value> <property>? Print each of them in SMILES format.”
“Can you generate fifty structurally diverse molecules with <value> <property>? Print each of them in SMILES format.”
Appendix C Additional results
Table 10: Visualization of the generated molecules from beam search. The $48$ molecules are generated with the description: “The molecule is a primary aliphatic ammonium ion which is obtained from streptothricin F by protonation of the guanidino and amino groups. It has a role as an antimicrobial agent. It is a guanidinium ion and a primary aliphatic ammonium ion. It is a conjugate acid of a streptothricin F’.”
Example of outputs (IntDiv. is $0.53$)
Table 11: Visualization of the generated molecules from diverse beam search. The $48$ molecules are generated with the description: “The molecule is a primary aliphatic ammonium ion which is obtained from streptothricin F by protonation of the guanidino and amino groups. It has a role as an antimicrobial agent. It is a guanidinium ion and a primary aliphatic ammonium ion. It is a conjugate acid of a streptothricin F’.”
Example of outputs (IntDiv. is $0.67$)
Table 12: Visualization of the generated molecules from constrastive beam search. The $48$ molecules are generated with the description: “The molecule is a primary aliphatic ammonium ion which is obtained from streptothricin F by protonation of the guanidino and amino groups. It has a role as an antimicrobial agent. It is a guanidinium ion and a primary aliphatic ammonium ion. It is a conjugate acid of a streptothricin F’.”
Example of outputs (IntDiv. is $0.58$)
Table 13: Visualization of the generated molecules from our method. The $48$ molecules are generated with the description: “The molecule is a primary aliphatic ammonium ion which is obtained from streptothricin F by protonation of the guanidino and amino groups. It has a role as an antimicrobial agent. It is a guanidinium ion and a primary aliphatic ammonium ion. It is a conjugate acid of a streptothricin F.”
Example of outputs (IntDiv. is $0.77$)